correct answers will be given two points, and compl. 1. Solve the equation by completing the square. (T x(2x + 5) = 3
Answer #1

# 20.1271 +5) = 3 2x² + 5x=3 2 = - 2x + 5x-3 = 0 2x +(632-x) -3 = 2 20 ²7 60-x-3=0 22(273)-1(x+3) = 0 (2x-1) -1)(x+3) = 0 22-Solution is:

x = -3

x = 1/2 Or 0.5